OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

jimthompson5910 (jim_thompson5910):

hint: \[\Large \text{sale price} = \text{purchase price}+0.40\times\left(\text{purchase price}\right)\] the "0.40" refers to the 40%. This is known as the markup.

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

let me know if that's enough to get started

OpenStudy (anonymous):

im still confused im not very good at algebra

jimthompson5910 (jim_thompson5910):

what's the purchase price?

OpenStudy (anonymous):

all i know is they need to make a $500 profit

jimthompson5910 (jim_thompson5910):

they give it to you in the problem

OpenStudy (anonymous):

i think they would need to sell 2500 flowers

jimthompson5910 (jim_thompson5910):

how did you get that number?

OpenStudy (anonymous):

0.20f = 500 f = 500/0.20

jimthompson5910 (jim_thompson5910):

very good, the sale price is $0.70 so they make 20 cents per flower

OpenStudy (anonymous):

ahh ok

OpenStudy (anonymous):

theres another question that is very hard for me